OpenStudy (lizz123):

MEDAL WILL BE GIVEN!! Write an equation in slope-intercept form the line passing through each pair of points. (9,-1) (6,-2)

Mehek (mehek14):

first find the slope

Mehek (mehek14):



OpenStudy (lizz123):

-2--1 -----> -2+1=-1 6-9=-3

Mehek (mehek14):

ok so your slope is -1/-3 which becomes 1/3 since both are positive so your equation is y = 1/3x + b now find b

OpenStudy (lizz123):

How do I find that I know how to find the slope but Not the b(y-inter) part though

Mehek (mehek14):

you subtract the two points y2-y1/x2-x1

OpenStudy (lizz123):

 Do I multiply this @Mehek14